A 1.0-L buffer solution contains 0.100 mol HC2H3O2 and 0.100 mol NaC2H3O2. The value of Ka for HC2H3O2 is 1.8×10−5.

Part A

Calculate the pH of the solution, upon addition of 0.095 mol of NaOH to the original buffer.

addition of 0.095 mol of NaOH to the original buffer.
Answer #1

after addition of NaOH following reaction takes place.

CH3COOH + NaOH --------> CH3COONa + H2O

[NaC2H3O2] = 0.100 + 0.095 = 0.195 M

[HC2H3O2] = 0.10 - 0.095 = 0.005 M

pH = pKa + log [NaC2H3O2] / [HC2H3O2]

pKa = - log Ka = - log [1.8 x 10-5] = 4.74

pH = 4.74 + log [0.195] / [0.005]

pH = 6.33